OpenStudy (anonymous):

I think I get it now :) Find a power series for f(x)=xln(1+x) \[f(x)=xln(1+x)\] \[\frac{d}{dx}xln(1+x)\] \[=\frac{x}{x+1}+ln(x+1)\] \[=\sum_{n=0}^{\infty}(-1)^nx^n+\frac{d}{dx}ln(x+1)\] \[\sum_{n=0}^{\infty}(-1)^nx^n+\sum_{n=0}^{\infty}(-1)^n\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

How did I do?

OpenStudy (anonymous):

ok since

OpenStudy (anonymous):

multiply the above series with x

OpenStudy (anonymous):

I keep making that error.

OpenStudy (anonymous):

Why is the \[ln(1+x)=\sum (-1)^{n-1}\frac{x^n}{n}\] and not \[\sum_{n=0}^{\infty}(-1)^n\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

I've worked on this problem before and I just keep getting it wrong. What's your thought process?

OpenStudy (anonymous):

ok let me do this completely so that you will understand it .

OpenStudy (rogue):

\[f (x) = \ln (1 + x) = \sum_{n = 1}^{\infty} (-1)^{n + 1} \frac {x^n}{n} = \sum_{n = 0}^{\infty} (-1)^n \frac {x^{n+1}}{n + 1}\]

OpenStudy (rogue):

The thing you get and the standard one are equivalent, yours is just shifted 1 back.

OpenStudy (anonymous):

i think they are the same right?

OpenStudy (rogue):

Yeah, they are.

OpenStudy (anonymous):

how did you do this shift

OpenStudy (rogue):

You can leave it with the shift, but it's just simpler to use the standard form and start the series from n = 1.

OpenStudy (anonymous):

Here is another one of my problems. when do I have n=0 and when do I have n=1?

OpenStudy (anonymous):

when n=0 doesn't yield a power series

OpenStudy (anonymous):

elaborate @Outkast3r09 Does that mean the last n=0 statement above is incorrect?

OpenStudy (anonymous):

no say if you get \[x^{-1}\]

OpenStudy (anonymous):

you'd move it up, or in this.. it's undefined at n=0

OpenStudy (anonymous):

ok so when were multiplying x to ln(x+1) how would that effect \[\sum_{n = 0}^{\infty} (-1)^n \frac {x^{n+1}}{n + 1}\] ?

OpenStudy (anonymous):

*we're

OpenStudy (anonymous):

it'd effect \[x^{n+1}\]

OpenStudy (anonymous):

it is simply just you are adding to the power of x

OpenStudy (anonymous):

\[x^nx^1x^1\]=\[x^{n+2}\]

OpenStudy (anonymous):

oh so... \[\sum_{n = 0}^{\infty} (-1)^n \frac {x^{n+2}}{n + 1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what if it's ln(x^2+1) that would make \[\int\frac{2x}{x^2+1}\]

OpenStudy (anonymous):

just to make clarification the series i found start at n=1 you started it at n=0 it is shifted at n=0

OpenStudy (anonymous):

\[\sum\int\frac{2x}{x^2+1}\] \[2\sum\int\frac{x}{x^2+1}\]

OpenStudy (anonymous):

it is different now.

OpenStudy (anonymous):

So I wouldn't start off by taking the derivative of ln(x^2+1) hmmmmm lets see

OpenStudy (anonymous):

you need to expand this in power series formula so you should have the form \[\Large \frac{1}{1-x}\]

OpenStudy (anonymous):

\[=2x\frac{1}{1-(-x^2)} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[2x\sum(-x^2)^n\] \[2x\sum(-1)^n(x^{2n})\]

OpenStudy (anonymous):

\[2\sum(-1)^nx^{3n}\]

OpenStudy (anonymous):

\[\sum(-2)^nx^{3n}\]

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

you here ?

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

so how did I do?

OpenStudy (anonymous):

could I write the sum from n=0 in this case? because that would make the first number a 1

OpenStudy (anonymous):

I did it wrong again. :(

OpenStudy (anonymous):

i hope you got it now :) and you can do any log series now :)

OpenStudy (anonymous):

my sum makes perfect sense to me.

OpenStudy (anonymous):

Is the idea that I replace x by x^2 in \[(-1)^n\frac{x^{n+1}}{n+1}\] which gives me \[(-1)^n\frac{(x^2)^{n+1}}{n+1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what happens to the 2 that was upfront?

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

THank you @sami-21 and everyone that helped!

OpenStudy (anonymous):

you are welcome:)

OpenStudy (anonymous):